# GeoSplat: A Deep Dive into Geometry-Constrained Gaussian Splatting

### v2

@@ 3.2 Noise-robust and Efficient Geometric Estimations
Previous works obtained noisy geometric priors through unreliable methods (e.g., local PCA), and the priors were also kept static during optimization. As Gaussian primitives got denser, those priors became even more unreliable. To address their limitations, we present noise-robust and efficient estimation methods that can provide dynamic geometric information.
@@
Our estimation method assumes that every Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ lies on a local manifold $\mathcal{M}$ (i.e., Definition 2.1). We summarize how it works as below, and due to the space limitation, we provide derivation details in Appendix E.
@@
We have explored a different geometric language: varifold. Similar to the manifold-based case, Buet et al. (2022) constructed a matrix: weak second fundamental form (WSFF), whose eigendecomposition provides full curvature information. In Appendix D, we adapt that matrix to our setting, with its expression $\mathcal{B}^{i}$ for a Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ as
@@
where $m_{j}$ is some weight, $\chi_{\epsilon},\Upsilon_{\epsilon}$ are kernel functions that depend on bandwidth $\epsilon$, $\mathrm{sym}(\square)=(\square+\square^{\top})/2$ is a symmetry operation for any matrix $\square$, and $\widetilde{\mathbf{V}}^{j}=\mathbf{I}-\widetilde{\mathbf{u}}^{i}_{3}(\widetilde{\mathbf{u}}^{i}_{3})^{T}$ is the projection matrix of tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$, with its basis as $\mathbf{V}^{i}=[\mathbf{u}^{i}_{1};\mathbf{u}^{i}_{2}]$.
@@ 4 Related Work
Common 3D scenes have a clear geometric structure (e.g., smoothness), so there is an emerging field in the literature (Wang et al., 2024; Bonilla et al., 2024; Li et al., 2025) that aims to regularize the optimization of Gaussian splatting in terms of certain geometric information. For example, Wang et al. (2024); Turkulainen et al. (2025) aligned the Gaussian primitive towards the normal direction, and Li et al. (2025) not only considered that, but also placed the densified primitives in the tangent plane. While achieving noticeable performance improvements, those earlier studies mainly considered low-order geometric quantities (e.g., normal vector), neglecting the valuable information contained in high-order ones (e.g., curvature). As a rare case, Ververas et al. (2025) adopted the mean curvature to identify flat areas, though this was not a fully correct strategy (as explained in Sec. 3.1.1) and their implementation was still based on low-order quantities. A major contribution of this paper is the introduction of GeoSplat: a geometry-constrained optimization framework that exploits the high-order geometric information neglected by prior works.
Previous studies were limited by unreliable estimation methods, which yielded static and noise-sensitive geometric priors. For example, Wang et al. (2024) adopted the StableNormal model (Ye et al., 2024) to predict normal vectors, which tends to fail on unseen data, while Li et al. (2025) employed Agglomerative Hierarchical Clustering (AHC) (Feng et al., 2014), a classical clustering method that is not robust to sparsity. Even worse, as Gaussian primitives become denser during optimization, these static priors become increasingly unreliable. Based on inherent geometric structures (e.g., local manifolds), we propose efficient estimation methods to provide dynamic and reliable geometric information during optimization.
@@ 5 Experiments
In this section, we show extensive experiment results that verify the effectiveness of our framework. We provide the details of our experiment setup and more results in Appendix L.
@@ 5.1 Main Results
We compare our models with a number of baselines (e.g., GeoGaussian) on two types of view synthesis datasets: Replica (Straub et al., 2019) and ICL (Handa et al., 2014), across $12$ different scenes. The results are provided in Table 1 and Table 2, showing that our optimization framework: GeoSplat, is able to significantly improve Gaussian splatting and outperform previous baselines. For the first point, we can see that our varifold-based model achieves consistent noticeable performance gains relative to 3DGS (Kerbl et al., 2023), which are $2.77\%$ on ICL Room-1 in terms of PSNR, and our manifold-based model also outperforms it by $0.93\%$ on ICL Office-2 in terms of SSIM. For the second point, we can see that both our manifold-based and varifold-based models significantly outperform the key baseline: GeoGaussian, on every Replica dataset. For example, we achieve higher PSNR scores by $3.42\%$ on Replica R1 and $2.01\%$ on Replica OFF4. The consistent performance gains over a dozen of datasets indicate that our framework is indeed effective.
@@ 6 Conclusion
In this work, we introduce a general framework of geometry-constrained Gaussian splatting, with an emphasize to exploit high-order geometric information (e.g., curvature) that is largely neglected by previous methods. The experiment results show that our framework significantly improves Gaussian splatting and outperforms previous baselines, with very few artifacts (e.g., outliers) in low-resource settings. Building on inherent geometric structures (e.g., local manifolds), we present efficient estimation methods that provide our framework with noise-robust and dynamic geometric priors, thereby overcoming the reliance on static and noise-sensitive priors in existing approaches.
@@ Appendix A Review of Basic Geometry
In this section, we will briefly review two different geometric structures: manifold (Lee, 2018) in differential geometry and varifold (Allard, 1972) in geometric measure theory.
@@ A.1 Manifold in Riemannian Geometry
A smooth manifold $\mathcal{M}$ in Riemannian geometry (Lee, 2018) is associated with Riemannian metric $g$. An informal understanding is to view it as a point set that locally looks like a vector space, with inner products defined variously for different local spaces. In the rigorous formulation, such a local space is conceptualized as the tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$ at every point $\mathbf{q}\in\mathcal{M}$, with metric $g_{\mathbf{q}}:\mathcal{T}_{\mathbf{q}}\mathcal{M}\times\mathcal{T}_{\mathbf{q}}\mathcal{M}\rightarrow\mathbb{R}$ as the inner product to measure both vector length $\|\mathbf{v}\|^{2}=g_{\mathbf{q}}(\mathbf{v},\mathbf{v}),\mathbf{v}\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$ and similarity $g_{\mathbf{q}}(\mathbf{v}_{1},\mathbf{v}_{2}),\mathbf{v}_{1},\mathbf{v}_{2}\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$. All tangent planes share the same dimension, which is consistent the defined dimension $D_{\mathcal{M}}\in\mathbb{N}$ of the global manifold, and their disjoint union: tangent bundle $\sqcup_{\mathbf{q}\in\mathcal{M}}\mathcal{T}_{\mathbf{q}}\mathcal{M}$, is denoted as $\mathcal{T}\mathcal{M}$. A smooth section $\mathbf{V}$ of the bundle (i.e., vector field on the manifold) is a map that pairs every point $\mathbf{q}$ with a tangent vector $\mathbf{V}({\mathbf{q}})\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$. We denote the space of all smooth sections as $\mathcal{V}(\mathcal{M})$, and represent the set containing all smooth function $f:\mathcal{M}\rightarrow\mathbb{R}$ as $\mathcal{C}^{\infty}(\mathcal{M})$. For the normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ at each point $\mathbf{q}$ on the manifold $\mathcal{M}$, it is easier to define it in a Euclidean ambient space $\mathcal{E}=\mathbb{R}^{D_{\mathcal{E}}}$ that isometrically embeds manifold $\mathcal{M}$ and has a larger dimension $D_{\mathcal{E}}>D_{\mathcal{M}}$. In that sense, the space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ consists of normal vector $\mathbf{n}\in\mathcal{E}$ that is perpendicular to the tangent space $\mathcal{T}_{\mathbf{q}}\mathcal{M}$.
@@
A.2 Another Language: Varifold
@@
While manifold is a popular geometric language, it has a strong assumption of the smooth structure. Varifold (Allard, 1972; Buet et al., 2017) is another mathematical tool free from this assumption. Given a countably rectifiable set (Federer, 2014) $\mathsf{M}$ that is contained in an open set $\Omega\subset\mathcal{E}$ and a non-negative function $\theta:\mathsf{M}\rightarrow\mathbb{R}$ that remains positive almost everywhere, the rectifiable varifold $\mathcal{W}$ is defined as a joint non-negative Radon measure $\theta\mathcal{H}^{D_{\mathsf{M}}}\otimes\delta_{\mathcal{T}_{\mathbf{\star}}\mathsf{M}}$ on the product space $\mathsf{M}\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$. Here $\mathcal{H}^{D_{\mathsf{M}}}$ denotes the $D_{\mathsf{M}}$-dimensional Hausdorff measure (Federer, 2014) for point set $\mathsf{M}$, $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ represents a Grassmannian manifold (Lee, 2018) formed by all $D_{\mathsf{M}}$-dimensional linear subspaces of the ambient space $\mathcal{E}$, and $\delta_{\mathcal{T}_{\mathbf{\star}}\mathsf{M}}$ represents a Dirac measure over the Grassmannian $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$, with its mass centering on the approximate tangent plane $\mathcal{T}_{\mathbf{\star}}\mathsf{M}$ of some point $\star\in\mathsf{M}$.
As a Radon measure, the varifold $\mathcal{W}$ can be generally formulated through integral:
@@
$\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\varphi(\mathbf{q},\mathsf{T})\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T})=\int_{\mathsf{M}}\varphi(\mathbf{q},\mathcal{T}_{\mathbf{q}}\mathsf{M})\theta(\mathbf{q})\mathrm{d}\mathcal{H}^{D_{\mathsf{M}}}(\mathbf{q}),$ (15)
where $\varphi$ is any continuous test function with compact support. Let $\mathsf{P}$ denote the natural projection from product space $\mathsf{M}\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ to the point set $\mathsf{M}$ as $\mathsf{P}(\mathbf{q},\mathsf{T})=\mathbf{q},\forall\mathsf{T}\in\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$, then the mass of varifold $\mathcal{W}$ on a Borel set $B$ can be expressed as
$\|\mathcal{W}\|(B)=\mathcal{W}(\mathsf{P}^{-1}(B))=\int_{B}\theta(\mathbf{q},\mathcal{T}_{\mathbf{q}}\mathsf{M})\mathrm{d}\mathcal{H}^{D_{\mathsf{M}}}(\mathbf{q}),$ (16)
where term $\mathsf{P}^{-1}(B)$ denotes the preimage.
A key concept in the varifold theory is the first-order variation $\vartheta\mathcal{W}$. By Riesz theorem (Yosida, 2012), the variation can be shaped as
$\int_{\Omega}\langle\vartheta\mathcal{W}(\mathbf{q}),\mathsf{V}(\mathbf{q})\rangle\mathrm{d}\mathcal{H}^{D_{\mathsf{M}}}(\mathbf{q})=\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\mathrm{div}^{\mathsf{T}}\mathsf{V}(\mathbf{q})\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T}),$ (17)
where $\mathsf{V}$ is any vector field supported on set $\mathsf{M}$, and $\mathrm{div}^{\mathsf{T}}$ denotes the divergence of field $\mathsf{V}$ with respect to a tangent plane $\mathsf{T}$. Importantly, according to Radon-Nikodym theorem (Halmos, 2013), the variation $\vartheta\mathcal{W}$ can be decomposed as
$\vartheta\mathcal{W}=-\mathsf{H}\|\mathcal{W}\|+\widetilde{\vartheta}\mathcal{W},$ (18)
where $\widetilde{\vartheta}\mathcal{W}$ is a singular vector-valued measure with respect to the mass $\|\mathcal{W}\|$ and $\mathsf{H}$ defines the generalized mean curvature.

### v3

@@ 3.2 Noise-robust and Efficient Geometric Estimations
Previous works obtained noisy geometric priors through potentially unreliable methods (e.g., local PCA), and the priors were also kept static during optimization. As Gaussian primitives got denser, those priors became even more unreliable. To address their limitations, we present noise-robust and efficient estimation methods that can provide dynamic geometric information.
@@
Our estimation method assumes that every Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ lies on a local manifold $\mathcal{M}$ (i.e., Definition 2.1). We summarize below how it works, and due to the space limitation, we provide derivation details in Appendix E.
@@
We have explored a different geometric approach based on varifolds. Generalizing the notion of second fundamental form valid for manifolds, Buet et al. (2022) introduced for point cloud varifolds an approximate second fundamental form (WSFF) as a matrix whose eigendecomposition provides full approximate curvature information. In Appendix D, we adapt that matrix to our setting, with its expression $\mathcal{B}^{i}$ for a Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ as
@@
where $m_{j}$ denotes a weight, $\chi_{\epsilon},\Upsilon_{\epsilon}$ are kernel functions that depend on the approximation scale $\epsilon$, $\mathrm{sym}(\square)=(\square+\square^{\top})/2$ is a symmetrization operation for any matrix $\square$, and $\widetilde{\mathbf{V}}^{j}=\mathbf{I}-\widetilde{\mathbf{u}}^{i}_{3}(\widetilde{\mathbf{u}}^{i}_{3})^{T}$ is the orthogonal projection matrix onto the tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$, with its basis as $\mathbf{V}^{i}=[\mathbf{u}^{i}_{1};\mathbf{u}^{i}_{2}]$.
@@ 4 Related Work
Common 3D scenes have a clear geometric structure (e.g., smoothness), so there is an emerging field in the literature (Wang et al., 2024; Bonilla et al., 2024; Li et al., 2025) that aims to regularize the optimization of Gaussian splatting in terms of certain geometric information. For example, Wang et al. (2024); Turkulainen et al. (2025) aligned the Gaussian primitive towards the normal direction, and Li et al. (2025) not only considered that, but also placed the densified primitives in the tangent plane. While achieving noticeable performance improvements, those earlier studies mainly considered low-order geometric quantities (e.g., normal vector), neglecting the valuable information contained in higher-order ones (e.g., curvature). As a rare case, Ververas et al. (2025) adopted the mean curvature to identify flat areas, though this was not a fully correct strategy (as explained in Sec. 3.1.1) and their implementation was still based on low-order quantities. A major contribution of this paper is the introduction of GeoSplat: a geometry-constrained optimization framework that exploits the higher-order geometric information neglected by prior works.
Previous studies were limited by potentially unreliable estimation methods, which yielded static and noise-sensitive geometric priors. For example, Wang et al. (2024) adopted the StableNormal model (Ye et al., 2024) to predict normal vectors, which tends to fail on unseen data, while Li et al. (2025) employed Agglomerative Hierarchical Clustering (AHC) (Feng et al., 2014), a classical clustering method that is not robust to sparsity. Even worse, as Gaussian primitives become denser during optimization, these static priors become increasingly unreliable. Based on inherent geometric structures (e.g., local manifolds), we propose efficient estimation methods to provide dynamic and reliable geometric information during optimization.
@@ 5 Experiments
In this section, we show extensive experiment results that verify the effectiveness of our framework. We provide the details of our experiment setup and more results in Appendix F.
@@ 5.1 Main Results
We compare our models with a number of baselines (e.g., GeoGaussian) on two types of view synthesis datasets, Replica (Straub et al., 2019) and ICL (Handa et al., 2014), across $12$ different scenes. The results are provided in Table 1 and Table 2, showing that our optimization framework, GeoSplat, is able to significantly improve Gaussian splatting and outperform previous baselines. For the first point, we can see that our varifold-based model achieves consistent noticeable performance gains relative to 3DGS (Kerbl et al., 2023), which are $2.77\%$ on ICL Room-1 in terms of PSNR, and our manifold-based model also outperforms it by $0.93\%$ on ICL Office-2 in terms of SSIM. For the second point, we can see that both our manifold-based and varifold-based models significantly outperform the key baseline, GeoGaussian, on every Replica dataset. For example, we achieve higher PSNR scores by $3.42\%$ on Replica R1 and $2.01\%$ on Replica OFF4. The consistent performance gains over a dozen of datasets indicate that our framework is indeed effective.
@@ 6 Conclusion
In this work, we introduce a general framework of geometry-constrained Gaussian splatting, with an emphasize to exploit higher-order geometric information (e.g., curvature) that is largely neglected by previous methods. The experiment results show that our framework significantly improves Gaussian splatting and outperforms previous baselines, with very few artifacts in low-resource settings. Building on inherent geometric structures (e.g., local manifolds), we present efficient estimation methods that provide our framework with noise-robust and dynamic geometric priors, thereby overcoming the reliance on static and noise-sensitive priors in existing approaches.
@@ Appendix A Review of Basic Geometry
In this section, we will briefly review two different geometric structures: manifolds (Lee, 2018) in differential geometry and varifolds (Allard, 1972; Menne, 2017) in geometric measure theory.
@@ A.1 Manifold in Riemannian Geometry
A smooth manifold $\mathcal{M}$ in Riemannian geometry (Lee, 2018) is associated with Riemannian metric $g$. An informal understanding is to view it as a point set that locally looks like a vector space, with inner products defined variously for different local spaces. In the rigorous formulation, such a local space is conceptualized as the tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$ at every point $\mathbf{q}\in\mathcal{M}$, with metric $g_{\mathbf{q}}:\mathcal{T}_{\mathbf{q}}\mathcal{M}\times\mathcal{T}_{\mathbf{q}}\mathcal{M}\rightarrow\mathbb{R}$ as the inner product to measure both vector length $\|\mathbf{v}\|^{2}=g_{\mathbf{q}}(\mathbf{v},\mathbf{v}),\mathbf{v}\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$ and similarity $g_{\mathbf{q}}(\mathbf{v}_{1},\mathbf{v}_{2}),\mathbf{v}_{1},\mathbf{v}_{2}\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$. All tangent planes share the same dimension, which is consistent with the defined dimension $D_{\mathcal{M}}\in\mathbb{N}$ of the global manifold, and their disjoint union, the tangent bundle $\sqcup_{\mathbf{q}\in\mathcal{M}}\mathcal{T}_{\mathbf{q}}\mathcal{M}$, is denoted as $\mathcal{T}\mathcal{M}$. A smooth section $\mathbf{V}$ of the bundle (i.e., vector field on the manifold) is a map that pairs every point $\mathbf{q}$ with a tangent vector $\mathbf{V}({\mathbf{q}})\in\mathcal{T}_{\mathbf{q}}\mathcal{M}$. We denote the space of all smooth sections as $\mathcal{V}(\mathcal{M})$, and represent the set containing all smooth function $f:\mathcal{M}\rightarrow\mathbb{R}$ as $\mathcal{C}^{\infty}(\mathcal{M})$. For the normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ at each point $\mathbf{q}$ on the manifold $\mathcal{M}$, it is easier to define it in a Euclidean ambient space $\mathcal{E}=\mathbb{R}^{D_{\mathcal{E}}}$ that isometrically embeds manifold $\mathcal{M}$ and has a larger dimension $D_{\mathcal{E}}>D_{\mathcal{M}}$. In that sense, the space $\mathcal{N}_{\mathbf{q}}\mathcal{M}$ consists of normal vector $\mathbf{n}\in\mathcal{E}$ that is perpendicular to the tangent space $\mathcal{T}_{\mathbf{q}}\mathcal{M}$.
@@
A.2 Another Tool: Varifolds
@@
While manifold is a popular geometric language, it relies on a strong assumption of smoothness. The notion of varifold (Allard, 1972; Simon, 1983; Menne, 2017) is another mathematical tool that requires only weak regularity. For the sake of clarity, we first focus on the simpler subclass of rectifiable varifolds. Given a countably $D_{\mathsf{M}}$-rectifiable set (Federer, 1969) $\mathsf{M}$ that is contained in an open set $\Omega\subset\mathcal{E}$ and a nonnegative function $\theta:\mathsf{M}\rightarrow\mathbb{R}$ that is positive $\mathcal{H}^{D_{\mathsf{M}}}$-almost everywhere, the rectifiable varifold $\mathcal{W}$ is defined as the nonnegative Radon measure $\theta\mathcal{H}^{D_{\mathsf{M}}}_{|\mathsf{M}}\otimes\delta_{\mathcal{T}_{x}\mathsf{M}}$ on the product space $\mathsf{M}\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$. Here $\mathcal{H}^{D_{\mathsf{M}}}_{|\mathsf{M}}$ denotes the $D_{\mathsf{M}}$-dimensional Hausdorff measure supported on $\mathsf{M}$ (Federer, 1969), $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ is the Grassmannian manifold (Lee, 2018) formed by all $D_{\mathsf{M}}$-dimensional linear subspaces of the ambient space $\mathcal{E}$, and $\delta_{\mathcal{T}_{x}\mathsf{M}}$ is the Dirac measure in $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ supported on the tangent plane $\mathcal{T}_{x}\mathsf{M}$ at $x\in\mathsf{M}$.
The action of $\mathcal{W}$ on a continuous function $\varphi$ compactly supported in $\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ is
@@
$\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\varphi(\mathbf{q},\mathsf{T})\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T})=\int_{\mathsf{M}}\varphi(\mathbf{q},\mathcal{T}_{\mathbf{q}}\mathsf{M})\theta(\mathbf{q})\mathrm{d}\mathcal{H}^{D_{\mathsf{M}}}(\mathbf{q}).$ (15)
Denoting as $\mathsf{P}$ the canonical projection from the product space $\mathsf{M}\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ onto $\mathsf{M}$, we define the mass $\|\mathcal{W}\|$ of $\mathcal{W}$ as the Radon measure such that, for any Borel set $B\subset\Omega$,
$\|\mathcal{W}\|(B)=\mathcal{W}(\mathsf{P}^{-1}(B))=\int_{\mathsf{M}\cap B}\theta(\mathbf{q})\mathrm{d}\mathcal{H}^{D_{\mathsf{M}}}(\mathbf{q}).$ (16)
A key concept in the theory of varifolds is the notion of first variation $\delta\mathcal{W}$ of a varifold $\mathcal{W}$: for any smooth compactly supported vector field $\mathsf{X}$
$\delta\mathcal{W}(\mathsf{X})=\int_{\Omega\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}}\mathrm{div}^{\mathsf{T}}\mathsf{X}(\mathbf{q})\mathrm{d}\mathcal{W}(\mathbf{q},\mathsf{T}),$ (17)
where $\mathrm{div}^{\mathsf{T}}$ denotes the tangential divergence with respect to $T$. If the linear functional $\delta\mathcal{W}$ is locally bounded, it follows from Riesz theorem (Yosida, 2012) and Radon-Nikodym theorem (Halmos, 2013) that it can be associated with a unique Radon measure that admits the decomposition
$\delta\mathcal{W}=-\mathsf{H}\|\mathcal{W}\|+{\delta}^{S}\mathcal{W},$ (18)
where $\delta^{S}\mathcal{W}$ is a vector-valued measure singular with respect to the mass $\|\mathcal{W}\|$, and $\mathsf{H}$ is called the generalized mean curvature. The latter coincides with the classical notion of mean curvature when $\mathcal{W}$ is a rectifiable varifold associated with a smooth set and a locally constant multiplicity. More generally, a generalized notion of second fundamental form can be defined for varifolds, see Buet et al. (2022). The first variation of a point varifold is not locally bounded in general, so the above decomposition does not hold. However, it is possible to define for point cloud varifolds consistent notions of approximate mean curvature Buet et al. (2017) and approximate weak second fundamental form, see Buet et al. (2022) and Appendix D.